The table below outlines sales data for machine equipment across four yearly quarters. Following the Seasonal Forecasting method, calculate the seasonal index for quarter 3 in year 2.

Quarter Year 1 Year 2
Quarter 1 100 120
Quarter 2 180 140
Quarter 3 190 195
Quarter 4 150 160
Total Total
620 615
Answer #1

Seasonal Index is the ratio of average value in a seaon to overall average value

Overall average sales = \frac{Sum of all sales}{no. of quarters} = \frac{100+180+190+150+120+140+195+160}{8}

Average Sales = \frac{1235}{8}= 154.375

Seasonal Average for Quarter 3 = \frac{Q3 sales Year 1 + Q3 Sales Year 2}{2} = \frac{190+195}{2} = \frac{385}{2}

Seasonal Average for Q3 = 192.5

Thus, Seasonal Index for Q3 Year 2 = \frac{192.5}{154.375} = 1.247

Thus Seasonality Index for Quarter 3 in year 2 is 1.247

Seasonality Index above 1 indicates the demand in that season is above average.